# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ III-D Switching Law
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0,\\ {\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u\left(t_{1}^{-}\right){\boldsymbol{b}}\right)=0.\end{dcases}$ (27)
Assume that $u\equiv u_{2}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ on ${\mathcal{S}}_{2}$. Then, $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0$ since ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$. Since $\exists\varepsilon>0$, s.t. $\forall t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$, it has $\left.\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)\leq 0$. Therefore, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\left(u_{2}-u\left(t_{1}^{-}\right)\right)\leq 0$. Since ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\not=0$ and $\left|u\left(t_{1}^{-}\right)\right|\leq\left|u_{2}\right|={u_{\mathrm{m}}}$, (22) holds.
Furthermore, consider the monotonicity of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}$ at $t_{1}$ to guarantee the feasibility of ${\mathcal{S}}_{2}$ where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}<0$ for $\left(t_{1},t_{1}+\varepsilon\right)$. Note that on ${\mathcal{S}}_{2}$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{2}{\boldsymbol{b}}$; hence, $\forall r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)$. Since $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, Lemma 2 implies that $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t. $\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall 1\leq r<\hat{r}_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$.
@@
$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall 1\leq r<\hat{r}_{p},\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (28)
Hence, $\exists\varepsilon>0$, $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. By Proposition 2, $\forall p\in\mathcal{P}_{1}$, $t\in\left[t_{0},t_{1}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$. So Proposition 3-3 holds. Similarly, Proposition 3-4 holds. ∎
@@
Proposition 3 provides a necessary and sufficient condition for the feasibility of adjacent arcs near the connection. Note that the equality constraints at the connection time, i.e., (12), (21), (23), and (25), are not independent since (12) implies the latter three equations for $1\leq r<r_{p}$. When examining feasibility near the connection time, it is sufficient to consider all inequality constraints and the maximal linearly independent set of the above equality constraints. Furthermore, Proposition 3 temporarily does not consider constraints on $u$ and ${\boldsymbol{x}}$, except for those in $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ which will be discussed in Section IV-A.
IV Keypoints for the Feasibility of the Optimal Trajectory
The switching law proposed in Section III can describe a given optimal trajectory and the feasibility near connections of arcs, but it fails to fully present information on the feasibility during an arc. To address this limitation, this section investigates the feasibility of the optimal trajectory on each arc, especially of the “keypoints” in an arc. Section IV-A analyzes some constraints of the end of an arc besides Proposition 3. The tangent marker is proposed in Section IV-B to describe the feasibility of the optimal trajectory tangent to the constrained boundary. Finally, Section IV-C proposes the augmented switching law and provides a sufficient and necessary condition for the feasibility near the keypoints.
@@ IV-A Additional Constraints at the End of An Arc
Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is a system behavior for $t\in\left[t_{0},t_{1}\right]$.
@@
      ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0$. $\exists 1\leq\hat{r}_{p}\leq n$, s.t. $\forall 1\leq r<\hat{r}_{p}$, ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)=0$, and ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)<0$.
    Similar conclusions hold for ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}^{-}\right)+d_{p}\leq 0$.
@@
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$. $\exists 1\leq\hat{r}_{p}\leq n$, s.t. $\forall 1\leq r<\hat{r}_{p}$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
@@
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$, and $\forall 1\leq r\leq n$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
    Similar conclusions hold for $\left|u\left(t_{0}^{+}\right)\right|,\left|u\left(t_{1}^{-}\right)\right|\leq{u_{\mathrm{m}}}$.
@@
Proposition 4-1 holds for the same reason as that of Proposition 3.
Consider Proposition 4-2. Proposition 2 implies that $u=-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}$. $\exists\varepsilon>0$, s.t. $u\left(t\right)\leq{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if (1) $\exists\varepsilon>0$, s.t. $u<{u_{\mathrm{m}}}$ for $\left(t_{0},t_{0}+\varepsilon\right)$, or (2) $\exists\varepsilon>0$, s.t. $u\left(t\right)\equiv{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$. By Proposition 2, (1) is equivalent to $\exists 0\leq\hat{r}_{p}\leq n$, s.t. $\forall 0\leq r<\hat{r}_{p}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, and $\frac{\mathrm{d}^{\hat{r}_{p}}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{\hat{r}_{p}}}\left(t_{0}\right)<0$. (2) is equivalent to $\forall 0\leq r\leq n$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$. Hence, Proposition 4-2 holds. ∎
@@
Proposition 4 encompasses discussions on the cases covered by Proposition 3, while Proposition 3 provides more details for $\mathcal{P}_{1}\cup\mathcal{P}_{2}$.
Based on Proposition 4, the additional end-constraint is defined as follows:
Definition 3.
Assume that ${\mathcal{S}}$ is a given arc for $t\in\left[t_{0},t_{1}\right]$. At an end of ${\mathcal{S}}$, i.e., $t_{i}$, $i\in\left\{0,1\right\}$, ${\mathcal{S}}$ and its adjacent arc, if it exists, correspond to the active constraints $\mathcal{P}$ and $\mathcal{P}^{\prime}$. $\mathcal{P}$ and $\mathcal{P}^{\prime}$ induce the equality constraints ${\boldsymbol{F}}{\boldsymbol{x}}\left(t_{i}\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ where ${\boldsymbol{F}}$ has full row rank. The additional end-constraint of ${\boldsymbol{x}}\left(t_{i}\right)$ is defined as ${\mathcal{E}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$. Among them, $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{i}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{i}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$ are induced by Proposition 4 for the active constraints of ${\mathcal{S}}$ and its adjacent arc at $t_{i}$, except for those in $\mathcal{P}\cup\mathcal{P}^{\prime}$. $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank.
Remark.
Given a feasible arc, the order $\hat{r}_{p}$ in Proposition 4 is directly determined. In Definition 3, the additional end-constraint ${\mathcal{E}}$ does not include the equality constraints induced by $\mathcal{P}\cup\mathcal{P}^{\prime}$. To avoid over-determination, the equality constraints in ${\mathcal{E}}$ and $\mathcal{P}\cup\mathcal{P}^{\prime}$ are made linearly independent through elimination. The elimination process is guaranteed by the feasibility of the given trajectory.

### v2

@@ III-D Switching Law
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0,\\ \forall r\in\left[r_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0,\\ {\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u\left(t_{1}^{-}\right){\boldsymbol{b}}\right)=0.\end{dcases}$ (26)
Assume that $u\equiv u_{2}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ on ${\mathcal{S}}_{2}$. Then, $\forall r\in\left[r_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0$ since ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$. Since $\exists\varepsilon>0$, s.t. $\forall t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$, it holds that $\left.\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)\leq 0$. Therefore, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\left(u_{2}-u\left(t_{1}^{-}\right)\right)\leq 0$. Since ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\not=0$ and $\left|u\left(t_{1}^{-}\right)\right|\leq\left|u_{2}\right|={u_{\mathrm{m}}}$, (21) holds.
Furthermore, consider the monotonicity of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}$ at $t_{1}$ to guarantee the feasibility of ${\mathcal{S}}_{2}$ where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}<0$ for $\left(t_{1},t_{1}+\varepsilon\right)$. Note that on ${\mathcal{S}}_{2}$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{2}{\boldsymbol{b}}$; hence, $\forall r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)$. Since $\forall r\in\left[r_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, Lemma 1 implies that $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t. $\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall r\in\left[\hat{r}_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$.
@@
$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (27)
Hence, $\exists\varepsilon>0$, $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. By Proposition 2, $\forall p\in\mathcal{P}_{1}$, $t\in\left[t_{0},t_{1}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$. So Proposition 4.3 holds. Similarly, Proposition 4.4 holds. ∎
@@
Proposition 4 provides a necessary and sufficient condition for the feasibility of adjacent arcs near the connection. Note that the equality constraints at the connection time, i.e., (11), (20), (22), and (24), are allowed linearly dependent. When examining feasibility near the connection time, it is sufficient to consider inequality constraints and all linearly independent equality constraints.
IV Keypoints for the Trajectory’s Feasibility
The switching law proposed in Section III can describe a given optimal trajectory and the feasibility near connections of arcs, while the feasibility during an arc remains inadequately investigated. To address this limitation, this section investigates the feasibility of the optimal trajectory on each arc, especially of keypoints in an arc. As formally defined in Definition 5, a keypoint, where the trajectory touches state constraints’ boundaries in isolation, directly influences the feasibility of the perturbed trajectory. For the example shown in Fig. 1, the time points at $\mathcal{T}_{1}^{(1)}$, $\mathcal{T}_{1}^{(2)}$, $\mathcal{T}_{2}^{(2)}$, and ends of arcs are keypoints.
Section IV-A analyzes additional constraints of an arc’s end besides Proposition 4. The tangent marker is proposed in Section IV-B to describe the feasibility of a BBS trajectory tangent to constraints’ boundaries. Finally, Section IV-C proposes the augmented switching law and provides a sufficient and necessary condition for the feasibility near the keypoints.
@@ IV-A Additional Constraints at the End of An Arc
Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is a system behavior for $t\in\left[t_{0},t_{1}\right]$.
@@
      ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0$. Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t. $\forall r\in\left[\hat{r}_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)=0$, and ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)<0$.
    Similar conclusions hold for $\left[t_{1}-\varepsilon,t_{1}\right]$.
@@
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$. Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t. $\forall r\in\left[\hat{r}_{p}-1\right]$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
@@
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$. Furthermore, $\forall r\in\left[n\right]$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
    Similar conclusions hold for $u\geq-{u_{\mathrm{m}}}$ and $\left[t_{1}-\varepsilon,t_{1}\right]$.
@@
Proposition 5.1 holds for the same reason as that of Proposition 4.
Consider Proposition 5.2. Proposition 2 implies that $u=-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}$. $\exists\varepsilon>0$, s.t. $u\left(t\right)\leq{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if (1) $\exists\varepsilon>0$, s.t. $u<{u_{\mathrm{m}}}$ for $\left(t_{0},t_{0}+\varepsilon\right)$, or (2) $\exists\varepsilon>0$, s.t. $u\left(t\right)\equiv{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$. By Proposition 2, (1) is equivalent to $\exists\hat{r}_{p}\in\left[0,n\right]\cap{\mathbb{N}}$, s.t. $\forall r\in\left[0,\hat{r}_{p}-1\right]\cap{\mathbb{N}}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, and $\frac{\mathrm{d}^{\hat{r}_{p}}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{\hat{r}_{p}}}\left(t_{0}\right)<0$. (2) is equivalent to $\forall r\in\left[0,n\right]\cap{\mathbb{N}}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$. Hence, Proposition 5.2 holds. ∎
${\mathcal{L}}=\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{(1)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{1}\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{(2)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{2}\dots\mathcal{E}_{N-1}\left({\mathcal{S}}_{N},\left\{{\mathcal{T}}_{k}^{(N)}\right\}_{k=1}^{M_{N}}\right)$ (28)
Consider two connected arcs ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$. Proposition 4 guarantees the feasibility of constraints active along ${\mathcal{S}}_{1}$ or ${\mathcal{S}}_{2}$. If some constraints are active at the ends of arcs but are inactive in ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ a.e., then additional constraints at the connection, as defined in Definition 3, should be introduced for feasibility based on Proposition 5.
Definition 3.
Consider two arcs ${\mathcal{S}}_{i}$ with active constraints $\mathcal{P}_{i}$ for $t\in\left[t_{i-1},t_{i}\right]$, $i\in\left\{1,2\right\}$. $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ induce the equality constraints ${\boldsymbol{F}}{\boldsymbol{x}}\left(t_{1}\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ where ${\boldsymbol{F}}$ has full row rank. By Proposition 5, state constraints except $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ and control constraints induce $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, where $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank. Then, the additional end-constraint of ${\boldsymbol{x}}\left(t_{1}\right)$ is defined as ${\mathcal{E}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$. Furthermore, if $\widehat{{\boldsymbol{C}}}$ and $\widehat{{\boldsymbol{F}}}$ are empty matrices, then ${\mathcal{E}}$ is called empty.
@@
Given a feasible arc, the order $\hat{r}_{p}$ in Proposition 5 is directly determined. In Definition 3, the additional end-constraint ${\mathcal{E}}$ does not include the equality constraints induced by $\mathcal{P}_{1}\cup\mathcal{P}_{2}$. To avoid over-determination, the equality constraints in ${\mathcal{E}}$ and $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ are made linearly independent through elimination. The elimination process is guaranteed by the feasibility of the given trajectory.